# CAPE: A CLIP-Aware Pointing Ensemble of Complementary Heatmap Cues for Embodied Reference Understanding

### v1

@@ 4.1 Results
Tab. 1 compares the performance of our method with existing approaches on the YouRefIt benchmark. Our model achieves SOTA results across all objects at IoU thresholds of 0.25 and 0.50, despite being trained for only 30 epochs on a single GPU with a batch size of 4. According to our experiments, this limited training setup affects bounding box precision, contributing to the suboptimal performance at , alongside limitations related to the backbone (see Grounding DINO in Tab. 1 for backbone). In contrast, Touch-Line [35] was trained for 200 epochs with a batch size of 56. This clearly demonstrates the efficiency and effectiveness of our approach. Tab. 2 presents a comparison between our method and the previous SOTA, Touch-Line, using two new metrics in this field: CLIP score and . The higher CLIP scores indicate that our predicted bounding boxes have stronger semantic alignment with the input text, suggesting better object identification and improved bounding box precision. Similarly, our method achieves the best performance under the metric, which quantifies the distance between the centers of the predicted and GT bounding boxes. This confirms that our model produces more spatially accurate predictions. In Fig. 3, we present samples images with predicted and GT bounding boxes. We compare our model with Touch-Line [35].
@@
| A | Baseline | 71.2 | 60.1 | 32.8 | 0.7409 | 0.2662 |
| B | A + object center prediction | 70.8 | 61.6 | 34.6 | 0.7339 | 0.2707 |
| C | A + W2F heatmap | 68.9 | 59.7 | 32.4 | 0.7353 | 0.2984 |
| D | A + H2F heatmap | 71.9 | 62.8 | 33.8 | 0.7375 | 0.2694 |
| E | B + W2F heatmap () | 69.6 | 60.7 | 31.5 | 0.7344 | 0.2770 |
| F | B + H2F heatmap () | 72.9 | 62.3 | 35.1 | 0.7373 | 0.2490 |
| G | A + W2F heatmap + H2F heatmap | 70.2 | 60.2 | 33.8 | 0.7345 | 0.2744 |
| H | Full model - Ensemble of E and F | 75.0 | 65.4 | 35.7 | 0.7983 | 0.2476 |
@@ 4.2 Analysis
In this section, we provide a detailed analysis of the optimality of our design choices, including heatmap generation and the ensemble method. Specifically, we examine various heatmap injection strategies and ensemble approaches, among other factors. The baseline in Tab. 3 (Setup A) refers to a straightforward training of our model without incorporating any of the proposed contributions.
@@
The object center prediction head and its associated loss yield improvements over the baseline in terms of mAP at and . However, they have a detrimental effect on mAP at , as well as on the CLIP Score and . Notably, when combined with the heatmap (Setup E and F in Tab. 3), we observe improvements across most metrics (see comparisons from Setup C to E and from Setup D to F).
@@
| Channel-wise input | 68.7 | 58.5 | 33.4 | 0.7348 | 0.2959 |
| Channel-wise feature | 72.6 | 58.4 | 27.9 | 0.7372 | 0.2598 |
| Embedding feature | 72.9 | 62.2 | 35.1 | 0.7373 | 0.2490 |
@@
We explore various methods for incorporating the heatmap into the model and present the results in Tab. 4. First, we examine the channel-wise input approach, which involves concatenating the heatmap with the input image and feeding a 4-channel input into the image encoder. However, this method leads to a slight drop in performance. This is likely because the image encoder struggles to extract meaningful supervision from the fourth channel, as the heatmap is sparsely represented compared to the image content. In the second approach, channel-wise feature, we adopt the head prompting strategy proposed in [51]. Specifically, we use an embedding layer to learn heatmap embeddings, which are then added to the visual tokens. While this method is computationally efficient and does not require retraining the image encoder, the resulting performance improvement for our task is minimal. Finally, in the embedding feature setup, we introduce an additional CNN-based encoder to process the heatmap. The resulting heatmap embedding is passed through a projection head and then concatenated with the image and text embeddings before being fed into the transformer encoder. This approach yields the most substantial performance improvement among the methods evaluated.
@@
| Confidence-Only | 73.4 | 64.1 | 34.1 | 0.7735 | 0.2470 |
| CLIP-Only (Top-1) | 73.5 | 64.2 | 35.4 | 0.7932 | 0.2391 |
| CLIP-Only (Top-2 + Threshold=0.95) | 73.8 | 64.1 | 35.5 | 0.7938 | 0.2377 |
| CLIP Fusion | 73.6 | 64.4 | 34.5 | 0.7379 | 0.2422 |
| CAPE | 75.0 | 65.4 | 35.7 | 0.7983 | 0.2476 |
@@
We evaluate several ensemble strategies to determine the final referent bounding box from the outputs of the two pointing models and present the results in Tab. 5. (1) Confidence-Only: selects the box with the highest model confidence score among Top-1 prediction of both models. (2) CLIP-Only (Top-1): chooses the box with the highest CLIP similarity score between the text and Top-1 predictions from both models. (3) CLIP-Only (Top-2 + Threshold): extends the previous method by also considering the top-2 predictions from each model, provided their confidence scores exceed . (4) CLIP Fusion: computes a normalized CLIP score and sums it with the model’s confidence score. The box with the highest total score is selected. This method considers the top-2 predictions from both models. (5) CAPE: a hybrid method that combines approaches (3) and (4). For small objects, defined by [11] as occupying less than 0.48% of the image area, CAPE uses method (4), since CLIP tends to perform unreliably on very small referents. For all other objects, it uses method (3). This adaptive strategy balances the strengths of both scoring signals across varying object sizes. Please note that the threshold used in method (3), as well as the decision to apply method (4) for small objects, were determined based on performance analysis on the validation set to avoid overfitting to the test set.
Finetuned vs. frozen text encoder
We investigate the impact of freezing the text encoder during training. While freezing the text encoder reduces computational cost, mitigates overfitting in low-resource settings, and stabilizes the training (see Appendix D), we observe that finetuning consistently yields better performance across almost all evaluation metrics (see Tab. 6). This suggests that adapting the text features to the specific task of embodied reference understanding provides a meaningful benefit.
| Text Encoder | IoU=0.25 | IoU=0.5 | IoU=0.75 | CLIP | |
| Frozen, Setup A | 69.8 | 59.5 | 33.1 | 0.7363 | 0.2994 |
| Finetuned, Setup A | 71.2 | 60.1 | 32.8 | 0.7409 | 0.2662 |
| Frozen, Setup F | 71.3 | 62.1 | 33.7 | 0.7382 | 0.2634 |
| Finetuned, Setup F | 72.9 | 62.2 | 35.1 | 0.7373 | 0.2490 |
How to create heatmap
We evaluate the Conic Attention Heatmap with and cone angles, as well as the Gaussian Ray Heatmap, which yields a more localized activation. Empirically, the Gaussian Ray Heatmap achieved the best performance (see Tab. 7). This suggests that broader attention regions may be less effective for the model, likely due to the high density of objects in typical scenes. Therefore, larger heatmaps tend to enclose multiple distractors, reducing the discriminative value of the spatial signal.
| Heatmap style | IoU=0.25 | IoU=0.5 | IoU=0.75 | CLIP | |
| Gaussian Ray Heatmap () | 72.9 | 62.2 | 35.1 | 0.7373 | 0.2490 |
| Conic Attention Heatmap () | 70.0 | 60.0 | 33.9 | 0.7392 | 0.2632 |
| Conic Attention Heatmap () | 71.5 | 59.4 | 30.8 | 0.7389 | 0.2764 |
@@ 5 Conclusion
We tackled the challenge of Embodied Reference Understanding by addressing the limitations of existing single-line pointing assumptions. Our proposed dual-model framework explicitly leverages both head-to-fingertip and wrist-to-fingertip lines, represented as Gaussian ray heatmaps, to provide richer supervision beyond following pointing line and improve pointing target detection. By integrating these complementary signals through our CLIP-Aware Pointing Ensemble (CAPE) module and enhancing the model with an auxiliary object center prediction task, we achieve notable improvements on the YouRefIt benchmark. Our results demonstrate that combining multiple pointing cues and multimodal signals leads to more accurate and robust referent understanding in complex visual scenes.
Limitations
Due to the lack of explicit depth information, the model occasionally selects an incorrect object that lies along the pointing direction but is not the intended referent. We plan to address this depth-related issue in future work. Additionally, the use of a CNN-based image backbone leads to suboptimal precision in bounding box predictions. This limitation can be mitigated by adopting a more advanced backbone architecture, such as the Swin Transformer.
@@
Appendix A Problem Analysis
How important is ERU task?
Embodied Reference Understanding (ERU) is a challenging multimodal task that combines visual, linguistic, and embodied cues (like gestures / pointing) to predict a target object in a scene. Despite recent progress, several key problems and open challenges remain. Natural language instructions are often vague or underspecified. Therefore, grounding models may detect multiple candidate objects or fail to resolve the correct referent altogether. Another problem is that when multiple instances of the same object class exist, models struggle to disambiguate. This results in over-detection (selecting all candidates) or under-confidence in selecting the target.
YouRefIt [11] introduced the ERU task, along with a multimodal dataset and a baseline approach to address it. However, the proposed method demonstrated limited performance, achieving a mean Average Precision (mAP) of only under an IoU threshold of . Subsequently, Touch-Line Transformer [35] proposed the concept of a Virtual Touch Line (VTL), based on the observation that the referent is typically aligned with the head-to-fingertip (H2F) line rather than the elbow-to-wrist (E2W) line. Since models struggle to follow pointing instructions, they propose using VTL prediction with a dedicated prediction head to encourage the model to focus more on the corresponding line. This approach led to a significant improvement in performance. However, our empirical analysis confirms that while the referent often aligns with the head-to-fingertip line, this is not always the case due to several factors:
If the person is looking elsewhere, such as at the camera, a robot, or another person, the alignment no longer holds.
The camera angle, consequently the image perspective, affects this alignment.
If the person is very close to the object or does not fully extend their arm while pointing the alignment may again be lost.
As a result, we find that in some cases, the referent aligns with the wrist-to-fingertip line. This indicates the need to incorporate both sources of information either in model training or decision step.
Appendix B Heatmap Generation
The heatmap generation method is a critical component of our approach. Recent work in this field [35] has shown that the referent often aligns with the head-to-fingertip line, referred to as the Virtual Touch-Line (VTL). Based on this insight, we generate a Gaussian Ray Heatmap using the head-to-fingertip line as a reference. While this assumption generally holds true, particularly when a person clearly points toward an object with an extended arm, it is not universally reliable. Variations occur both in this benchmark and in real-world scenarios. For instance, the alignment may be disrupted if the person is looking elsewhere (e.g., at a robot, camera, or another person). Similarly, camera perspective can also distort the alignment. In some cases, we observe that the wrist-to-fingertip line aligns more accurately with the referent than the head-to-fingertip line. To address this variability, we leverage both types of information and generate separate heatmaps for each sample accordingly.
@@ Appendix B Heatmap Generation
B.1 Head-to-Fingertip Heatmap
To generate this heatmap, we use the head and fingertip as reference points to draw a line starting from the head and extending toward the edge of the image in the direction of the fingertip. In the YouRefIt dataset [11], we have access to detailed annotations from [11, 35], which include coordinates for the eyes, elbows, wrists, and fingertips. We utilize these annotations to construct the reference line and generate the corresponding heatmap. In real-world applications where such annotations are not available, estimating these key points is straightforward. For the eye position, face detection followed by selecting the center point of the detected face yields comparable performance. Likewise, for wrist and fingertip locations, modern pose estimation models provide reliable predictions.
B.2 Wrist-to-Fingertip Heatmap
We follow the same pipeline used for generating the head-to-fingertip heatmap, but replace the eye coordinates with wrist coordinates and the fingertip coordinates remain the same.
B.3 Merging
For the unimodal experiments, we merge the head-to-fingertip and wrist-to-fingertip heatmaps to combine information from different pointing directions. However, this approach yields lower accuracy compared to using two separate models. We hypothesize that this is primarily because the model cannot inherently determine which pointing line is more appropriate for each sample.
B.4 Heatmap Injection Methods
We investigate three different methods for injecting heatmaps and conduct experiments using the head-to-fingertip heatmap.
Channel-wise input
In this method, we concatenate the input image with the heatmap and feed it into the image encoder, resulting in a 4-channel input. Consequently, this approach is not compatible with pretrained image encoders unless they are finetuned, as it requires adapting the model to handle the additional channel. Experimental results show that this method yields the least accurate performance.
Channel-wise feature
In this method, we follow a recently proposed approach [51] that has shown strong performance in the gaze target estimation literature. Specifically, we use a single learnable embedding layer to obtain embeddings of the heatmap, which are then added to the image embeddings.
Embedding feature
In this approach, we propose using an additional encoder. Given the sparse nature of the heatmap representation, we select a simpler encoder than the main image encoder. In this work, we employ ResNet-18 as the heatmap encoder. After obtaining the embedded representation, we pass it through a projection layer to match the required dimensionality. We then concatenate it with the image and text embeddings and feed the combined representation into the transformer encoder.
Appendix C Ensemble Methods
@@ C.2 CLIP-Only (Top-1)
C.3 CLIP-Only (Top-2 + Threshold)
@@
This method is similar to CLIP-Only (Top-1), where we take predictions from both models and compute the CLIP similarity score between the cropped bounding boxes and the input text. However, unlike the Top-1 variant, we also consider the Top-2 predictions from each model. However, we do not always include the second predictions; instead, we apply a confidence threshold determined empirically on the validation set, . If a model’s second-highest prediction has a confidence score greater than or equal to this threshold, it is included in the comparison. Finally, we select the bounding box with the highest CLIP similarity score as the final prediction. The algorithm for this method is presented in Algorithm C.3.
C.4 CLIP Fusion

### v4

@@ 4.1 Results
Tab. 1 compares the performance of our method with existing approaches on the YouRefIt benchmark. Our model achieves SOTA results in most cases, despite being trained for only 30 epochs on a single GPU with a batch size of 4. In contrast, Touch-Line models [38] were trained for 200 epochs with a batch size of 56. This clearly demonstrates the efficiency and effectiveness of our approach. However, our optimized training setup affects bounding box precision, contributing to the suboptimal performance at , alongside limitations related to the simpler backbone. At , Grounding DINO with a more complex backbone achieves the best performance.
Further, we evaluate recent LMMs on this task, and they fall significantly short of our model’s performance, despite being among the largest models and trained for visual grounding purposes. These results show using complementary pointing heatmap modalities in addition to image and text is essential. Besides, despite having two separate models, our approach shows better inference-time performance (2.44 fps) compared to Paligemma2 (1.24fps) and Qwen2.5vl (0.5fps). Tab. 2 presents further comparison using two new metrics in this field: CLIP score and . Higher CLIP scores indicate that our predicted bounding boxes have stronger semantic alignment with the input text, suggesting better object identification and improved bounding box precision. Similarly, our method achieves the best performance under the metric. This confirms that our model produces more spatially accurate predictions. In Tab. 3, the experimental results on unseen ISL dataset show that our CAPE model surpasses the SOTA Touch-Line model as well as the evaluated LMMs, except for the IoU=0.75 setup. This demonstrates the strong generalization capability of our method. In Fig. 3, we present sample images with predicted and GT bounding boxes. It is clearly seen from LMMs’ predictions, ambiguous text without pointing cues results in poor object localization outside of the pointing area and confusion with similar objects. For instance, Paligemma2 in third example and Qwen2.5vl in second and fourth examples detect wrong objects. On the other hand, the LMMs are strong at providing precise bounding box predictions due to extensive training and a robust backbone. Furthermore, even if Touch-Line model use pointing cues since it follows a pointing line instead of heatmaps, this let model to detect other objects in the pointing direction. In contrast, as shown in the examples, CAPE leverages explicit pointing heatmaps, enabling it to more accurately localize the intended referent. Also, it enables to choose final prediction from which of the more accurate model.
@@
| A | Baseline | 71.2 | 60.1 | 32.8 | 0.2469 | 0.2662 |
| B | A + object center prediction | 70.8 | 61.6 | 34.6 | 0.2446 | 0.2707 |
| C | A + W2F heatmap | 68.9 | 59.7 | 32.4 | 0.2451 | 0.2984 |
| D | A + H2F heatmap | 71.9 | 62.8 | 33.8 | 0.2458 | 0.2694 |
| E | B + W2F heatmap () | 69.6 | 60.7 | 31.5 | 0.2448 | 0.2770 |
| F | B + H2F heatmap () | 72.9 | 62.3 | 35.1 | 0.2457 | 0.2490 |
| G | A + W2F heatmap + H2F heatmap | 70.2 | 60.2 | 33.8 | 0.2448 | 0.2744 |
| H | Full model - Ensemble of E and F | 75.0 | 65.4 | 35.7 | 0.2661 | 0.2476 |
@@ 4.2 Analysis
In this section, we provide a detailed analysis of the optimality of our design choices, including heatmap generation and the ensemble method. The baseline in Tab. 4 (Setup A) refers to a straightforward training of our model without incorporating any of the proposed contributions.
@@
The object center prediction improves mAP at and but reduces performance at , as well as CLIP Score and . Notably, when combined with heatmap (Setup E and F in Tab. 4), it boosts most metrics. The object center prediction alone enhances geometric precision more than semantic alignment, while pairing it with spatially dense cues makes it a more effective refining signal.
@@
| Channel-wise input | 68.7 | 58.5 | 33.4 | 0.2449 | 0.2959 |
| Channel-wise feature | 72.6 | 58.4 | 27.9 | 0.2457 | 0.2598 |
| Embedding feature | 72.9 | 62.2 | 35.1 | 0.2458 | 0.2490 |
@@
We compare several heatmap injection strategies in Tab. 5. The channel-wise input method, which feeds the heatmap as a fourth image channel, slightly hurts performance, likely because the heatmap is sparsely represented compared to the image content. Inspired by the head location prompting strategy in [52], the channel-wise feature approach adds learned heatmap embeddings to the visual tokens, which is lightweight but provides only limited gains. In contrast, the embedding feature setup uses a dedicated CNN to encode the heatmap before fusing it with image and text embeddings, and delivers the strongest improvements across all metrics. The experiments show that effective heatmap injection requires the model to interpret the heatmap as a structured spatial signal rather than as a raw image channel or a light additive bias on visual tokens. Besides, the improvements seen in the embedding feature method support the gains observed in the main ablations (Setups D and F), where the H2F heatmap consistently strengthens localization and reduces .
@@
| Confidence-Only | 73.4 | 64.1 | 34.1 | 0.2578 | 0.2470 |
| CLIP-Only (Top-1) | 73.5 | 64.2 | 35.4 | 0.2644 | 0.2391 |
| CLIP-Only (Top-2 + Threshold=0.95) | 73.8 | 64.1 | 35.5 | 0.2646 | 0.2377 |
| CLIP Fusion | 73.6 | 64.4 | 34.5 | 0.2459 | 0.2422 |
| CAPE | 75.0 | 65.4 | 35.7 | 0.2661 | 0.2476 |
@@
We evaluate several ensemble strategies for selecting the final referent box from the two pointing models, as shown in Tab. 6. (1) Confidence-Only selects the Top-1 prediction with the highest model confidence. (2) CLIP-Only (Top-1) chooses the Top-1 box with the highest CLIP similarity. (3) CLIP-Only (Top-2 + Threshold) additionally considers each model’s top-2 predictions when their confidence scores exceed . (4) CLIP Fusion combines normalized CLIP and confidence scores from the top-2 predictions of both models and selects the box with the highest total score. (5) CAPE adaptively switches between (3) and (4): it uses (4) for small objects, where CLIP alone is unreliable, and (3) otherwise. The confidence threshold in (3) and the small-object rule in CAPE were tuned on the validation set to avoid test-time bias. Across these strategies, the hybrid design in CAPE aligns with observations from the main ablations: the two models provide complementary strengths, CLIP-based scoring refines semantic alignment for most objects, and confidence-based cues are especially important for small or ambiguous targets. This contributes to the strong gains seen in the full model (Setup H).
Failures
In Fig. 4, we present examples of failure cases. In the first example (column), although our model correctly detects the target object, CAPE selects the prediction from , which is incorrect in this case. In the second example, both of our models, as well as the Touch-Line models, produce incorrect predictions. However, guided by the pointing information, the predictions from and lie on the pointing line and are very close to the GT object. In contrast, the Touch-Line predictions are entirely off the pointing line and significantly distant from the GT object. In the third column, the bottles are extremely difficult to distinguish, even with explicit pointing cues. Nevertheless, the predicted bounding boxes cover other bottles adjacent to the GT one, which is acceptable, as they overlap with the pointing line and are semantically consistent with the text.
@@ 5 Conclusion
We address Embodied Reference Understanding by overcoming the limitations of relying on a single pointing line, using our dual-model framework to leverage complementary Gaussian ray heatmaps for richer supervision and improved target detection. Combined through the CLIP-Aware Pointing Ensemble (CAPE) and enhancing the model with an auxiliary object center prediction, our approach achieves strong improvements on the YouRefIt, ISL pointing, and CAESAR datasets. Our results demonstrate that combining multiple pointing cues and in general, multimodal signals, leads to more accurate and robust referent understanding in complex visual scenes. Beyond empirical gains, our work demonstrates a general principle of multimodal cue integration: fusing complementary spatial and semantic signals resolves ambiguities that single modalities cannot, providing a principle for robust referent understanding in complex visual scenes.
Limitations and Future Work
The model occasionally selects incorrect objects along the pointing direction due to the absence of depth information, and the CNN backbone limits bounding box precision. Future work could incorporate depth or 3D cues and adopt advanced backbones such as Swin Transformers. Additionally, formalizing a theoretical framework for multimodal cue integration and exploring adaptive weighting or attention mechanisms could further improve flexibility and robustness, allowing the system to optimally leverage multiple cues across diverse scenarios.
Acknowledgment
This work was supported in part by the European Union’s Horizon research and innovation program, project Meetween (101135798) and project DVPS (Diversibus Viis Plurima Solvo) (101213369), and KIT Campus Transfer GmbH (KCT) under contract from Carnegie-AI LLC.
@@ Acknowledgment
Appendix A Heatmap Generation
The heatmap generation method is a key component of our approach. Prior work [38] shows that the referent often aligns with the head-to-fingertip line, which inspired our Gaussian Ray Heatmap generation. While this generally holds when a person points clearly with an extended arm, it is not always reliable. Alignment can be disrupted if the person looks elsewhere (e.g., at a robot, camera, or another person), due to camera perspective, or when a person is close to the object and points using wrist motion. In such cases, the wrist-to-fingertip line aligns more accurately with the referent. To handle this variability, we generate separate heatmaps for both cues.
@@ Appendix A Heatmap Generation
A.1 Head-to-Fingertip Heatmap
To generate this heatmap, we take the head and fingertip as reference points and construct a ray starting at the head and extending toward the image boundary in the fingertip direction. In the YouRefIt dataset [13], we use the detailed annotations provided in [13, 38], which include eye, elbow, wrist, and fingertip coordinates. These keypoints allow us to reliably define the reference direction for heatmap generation.
In real-world settings where such annotations are unavailable, estimating these points is straightforward. For eye position, face detection followed by using the center of the detected face provides a strong approximation. Similarly, modern pose estimation models offer accurate wrist and fingertip predictions. For our experiments on the ISL Pointing and CAESAR datasets, we obtain necessary coordinates using OpenPose [9], which delivers robust and reliable keypoint detection.
A.2 Wrist-to-Fingertip Heatmap
We follow the same pipeline used for head-to-fingertip heatmap generation, but replace the eye coordinates with wrist coordinates and the fingertip coordinates remain the same.
A.3 How to Create Heatmap
We evaluate the Conic Attention Heatmap using and cone angles, as well as the Gaussian Ray Heatmap, which produces more localized activation. As shown in Tab. A.7, the Gaussian Ray Heatmap achieves the best performance. This suggests that broader attention regions may be less effective in practice, likely due to the high density of objects in typical scenes (see Fig. A.5). Therefore, larger heatmaps tend to cover multiple distractors, reducing the discriminative value of the spatial signal.
| Heatmap style | IoU=0.25 | IoU=0.5 | IoU=0.75 | CLIP | |
| Gaussian Ray Heatmap () | 72.9 | 62.2 | 35.1 | 0.2457 | 0.2490 |
| Conic Attention Heatmap () | 70.0 | 60.0 | 33.9 | 0.2464 | 0.2632 |
| Conic Attention Heatmap () | 71.5 | 59.4 | 30.8 | 0.2463 | 0.2764 |
Appendix B Ensemble Methods
@@ B.2 CLIP-Only (Top-1)
B.3 CLIP-Only (Top-2 + Threshold)
@@
This method is similar to CLIP-Only (Top-1), where we take predictions from both models and compute the CLIP similarity score between the cropped bounding boxes and the input text. However, unlike the Top-1 variant, we also consider the Top-2 predictions from each model. However, we do not always include the second predictions; instead, we apply a confidence threshold determined empirically on the validation set, . If a model’s second-highest prediction has a confidence score greater than or equal to this threshold, it is included in the comparison. Finally, we select the bounding box with the highest CLIP similarity score as the final prediction. The algorithm for this method is presented in Algorithm B.3.
B.4 CLIP Fusion